OpenStudy (sissyedgar):

The value of y varies directly with x², and y = 64 when x = 4. What is the value of y when x = 6? |___|

OpenStudy (sissyedgar):

@ShelbyWyatt

OpenStudy (usukidoll):

are you sure it's not cubed? I'm picturing this y = x^2 and then to get y = 64 I have to have x = 8...maybe it's just me tired from exams

OpenStudy (lncognlto):

its y = k(x^2)

OpenStudy (lncognlto):

:P how are exams going? @UsukiDoll

OpenStudy (sissyedgar):

The x is cubed... x² And which number is the K and which the x?

OpenStudy (usukidoll):

just finished my last one... I feel the same way as yesterday.

OpenStudy (sissyedgar):

Wait.. is it 8?

OpenStudy (usukidoll):

is it possible to show a screenshot of the problem so I can see if x^3 or x^2?

OpenStudy (sissyedgar):

It's x^2... But here's the screenshot...

OpenStudy (lncognlto):

well congrats! mine start next week so I bid you adieu for now.

OpenStudy (usukidoll):

wow... it could be x^2 k = 64 if x = 4 4^2k=64 16k=64 k = 64/16 = 4 so maybe now we need to find y = 6^2 (k) ?

OpenStudy (shelbywyatt):

y=(x*x)*4

OpenStudy (usukidoll):

x = 6 was mentioned so y = 6 x 6 and from earlier that k was 4 so our real equation to solve is y = 6 x 6 x 4

OpenStudy (sissyedgar):

Soooo.... y=(6*4)*4?

OpenStudy (shelbywyatt):

6*6=36 36*4

OpenStudy (sissyedgar):

36 * 4 = 144

OpenStudy (shelbywyatt):

yeah thats my answer

OpenStudy (sissyedgar):

Okay. Thankz guyz!!